OpenStudy (anonymous):

Martha has $180. She needs a total of $2,000 to start investing in stocks. She earns $60 per day working, of which she saves $50. Write an equation that Martha can use to determine the number of days, d, she needs to work to reach her goal of $2,000?

OpenStudy (anonymous):

2000/50=d = 40 days

OpenStudy (anonymous):

wait hold on.... 2000-180 already saved /50 =d

OpenStudy (anonymous):

are u sure this is right because i only have 2 attempts to take this quiz and really don't want to get it wrong & fail it again

OpenStudy (anonymous):

so 2000-180=$1820 left to save divided by $50 savings per day = d

OpenStudy (anonymous):

yes I'm sure

OpenStudy (anonymous):

so whats the exact equation ?

OpenStudy (anonymous):

2000-180=1820/50=d

OpenStudy (anonymous):

ok i'll check if its right or wrong